# Addressing the gravitational collapse of a massless scalar field with Physics-Informed Neural Networks

### v1

@@ Abstract
The gravitational collapse of a massless scalar field remains a demanding benchmark for numerical methods in numerical relativity, as it exhibits critical behavior at the boundary between dispersion and black hole formation. In this work we revisit this problem by relying on Physics-Informed Neural Networks (PINNs) as flexible partial differential equations solvers thereby providing a comparative assessment of several recent neural architectures. Building on the Einstein-massless-Klein-Gordon formulation in polar-areal coordinates, we consider four initial-value problems encompassing subcritical, critical, and supercritical regimes and use high-resolution finite-difference simulations as reference solutions. Our study is primarily comparative: we evaluate several state-of-the-art deep learning architectures, including vanilla and high-precision PINNs, sinusoidal-feature and quadratic-residual variants, Kolmogorov-Arnold Networks, all trained under a common loss design that encodes the field equations, boundary conditions, and causal time-space enforcement, together with a novel adaptive spacetime sampling. Within this framework we also introduce ModPINN, a modest modification of standard PINNs that augments standard multiplayer perceptrons with coordinate embeddings, quadratic layers among other common ingredients in recent literature. This study shows that deep-learning-based methods can reproduce finite-difference solutions for the scalar field and the spacetime metric with competitive accuracy using significantly fewer collocation points than more traditional methodologies. While no single architecture dominates in all regimes, ModPINN achieves particularly stable and accurate solutions near criticality, indicating that suitable designed embeddings and adaptive sampling can enhance the robustness of PINNs for challenging gravitational-collapse scenarios.
@@ 1.1 Problem setup and basic equations
The study of gravitational collapse and black-hole formation stands as one of the most interesting and revealing problems in modern physics. Beyond its astrophysical relevance, gravitational collapse provides an exceptional framework for testing the nonlinear, strong-field dynamics of general relativity (GR) and fundamental conjectures such as cosmic censorship. A particularly insightful and motivating aspect of this field arises in the so-called critical collapse problem, first discovered by Choptuik in the early 1990s [Choptuik_seminal], where the threshold between dispersion and black hole formation in the collapse of a massless scalar field was found to display a universal and self-similar behavior.
@@
The parameter $p$ corresponds to a specific characteristic of the initial distribution of the scalar field, $\phi(t=0,r):=\phi_{0}(r)$, where $t$ and $r$ denote the time and radial components, respectively, in polar-areal coordinates, $\{t,r,\varphi,\theta\}$. These coordinates help to write the time-dependent Schwarzschild-like metric
@@
The parameter $p$ naturally divides the entire set of initial conditions into two distinct categories: those that lead to the formation of a black hole during evolution and those that do not. Therefore, we can introduce the concept of a “threshold” when discussing the formation of an event horizon. Spacetimes evolving from initial conditions near this threshold exhibit properties that, from a mathematical perspective, bear an analogy to a critical phase transition in statistical mechanics (interested readers are referred to Sections 2.2 and 3.3 in [Gundlach2007]). In this manner, the parameter $p$ presents a critical value $p^{*}$ separating “sub-critical” solutions for which the mass dissipates ($p<p^{*}$) from “super-critical” solutions for which an event horizon forms ($p>p^{*}$), leading to the emergence of an apparent horizon.
@@
The set $(\Phi,\Pi)(t,r)$ in Eq. (8) is directly associated with the spatial and temporal derivatives of the scalar field, respectively. In terms of these variables, the set $(X,Y)(t,r)$ in Eq. (9) can be defined, leading to some relevant mathematical relations involving the derivatives of the mass aspect function, $m(t,r)$,
@@
For completeness, the spacetime curvature scalar $R$ can also be written in terms of this pair of variables as,
@@
Within this formulation, this parameter satisfies $C\in[0,1)$, where the limit $C\to 1$ characterizes the onset of gravitational collapse. From a computational perspective, the formation of an apparent horizon in numerical simulations would be identified by imposing a threshold $\epsilon$ such that $1-C\leq\epsilon$. In this manner, classical numerical methods based on finite-difference methods (FDM) can effectively solve the field equations in both dispersive scenarios of $\phi(t,r)$ and cases involving gravitational collapse (e.g. see [Choptuik_1998]). Despite the coupling between the different variables, the dynamical evolution is governed by Eq. (5), while Eqs. (6) and (7) serve as metric constraints.
@@ 1.2 Deep-Learning-based methodologies
Despite the many numerical studies, it is worth noticing that the most common methodology to solve the equations has been based on FDM. Only recent works have considered methods alternative to FDM to investigate gravitational collapse. In particular, the work of [finzi_2023] introduced a neural-based solver that mitigates artificial neural network (ANN) ill-conditioning while exhibiting linear scalability with respect to the number of parameters. Other studies, such as [hatefi_2023], developed a model also based on ANNs to investigate critical phenomena in Einstein-axion-dilaton models, finding no evidence of self-similar behavior. On the other hand, [hatefi_2023_2] and [hatefi_2024] employ Bayesian deep learning models to study the axion-dilaton system, utilizing Sequential Monte Carlo (SMC) sampling and related algorithms to obtain direct estimates of the critical exponent $\gamma$.
All of these recent studies establish a foundation for the incorporation of deep learning (DL) techniques in the investigation of gravitational collapse scenarios. The present work proposes the adoption of novel methodologies based on ANNs tailored to solve systems of partial differential equations (PDEs) and extract their underlying physics. Since their initial appearance in the scientific literature [seminal_pinns], Physics-Informed Neural Networks (PINNs) have emerged as a robust alternative to solve numerical problems governed by PDEs. These methods leverage the universal approximation theorem of ANNs [Pinkus_1999] to minimize the governing physical equations, incorporating them into the loss function used to train the network parameters. This function is typically augmented with physical constraints—commonly referred to as “inductive biases”—such as initial and boundary conditions, additional physical information to be minimized, or even agreement with experimental data. Therefore, our primary objective will be to demonstrate that the PINN methodology is fully capable of solving the EMKG equations (5-6), effectively addressing both the field dispersion and the gravitational collapse scenarios. This approach is proposed as a viable alternative to traditional numerical methods.
The paper is structured as follows. Section 2 presents an overview of recent literature on PINNs, to motivate our study. Next, Section 3 introduces our PINN-based methodology, which consists of a modified version of the standard (vanilla) version; this section also details the specific formulation used to minimize the loss function. Section 4 presents the results obtained for various initial values of the scalar field, including the critical case, with particular emphasis on evaluating the predictive performance of the model in that regime. Finally, Section 5 presents our conclusions and outlines potential directions for future research.
@@ 2 Review of recent literature
In general terms, the choice of neural architecture is a critical factor, as the number of trainable parameters must be appropriate for the problem at hand, whatever the specific task may be—whether solving PDEs or addressing another application. Among the aspects that define a neural architecture, several may be highlighted: the activation functions employed; the number of layers (depth) and their size (width); whether residual or skip connections are incorporated between layers; and whether any form of embedding is used, among other considerations. Several studies [Yifan_2024] conclude, in a range of independent physical problems, that relatively shallow but wide networks often outperform deeper but narrower ones. This reveals a concern that usually appears throughout the literature on multilayer perceptrons (MLPs) applied to physics, motivating more recent developments such as Physics-Informed Residual AdapTiVe Networks (PirateNets) [piratenets_2024], which are introduced as residual architectures that allow the model to start from shallow networks, progressively “revealing” additional layers only when they are advantageous to reduce the objective cost function. These networks also employ embeddings of the physical coordinates, mapping them into a high-dimensional feature space using random Fourier features [Tancik_2020]. However, although these embeddings can mitigate the frequency bias typically observed in PINNs [Sifan_2021] and facilitate the solution of physical problems involving high-frequency solutions, the selection of the maximum frequency scale and the standard deviation of the Gaussian distribution used to sample the frequencies are critical; if chosen poorly, these become problem-dependent hyperparameters that can hinder performance. Other architectures, such as the sf-PINNs proposed in [Wong_2024], however, apply a mapping of the input coordinates to sinusoidal features, encoding them after the first layer, thus increasing the variability of the input gradient and also helping mitigate the said frequency bias.
@@
Stepping somewhat away from traditional architectures, methodological variants such as the Quadratic Residual Networks (QRes) [qres_2021] aim to replace the standard linear layer with an alternative that introduces second-order interactions between the trainable parameters and the input, becoming more parameter-efficient and faster to train, in general, on scientific problems. Within the most recent DL literature, Kolmogorov-Arnold Networks (KANs) [kan_2025] have emerged as an alternative to conventional MLPs which have been also applied to PINNs. These mainly consist of substituting the trainable weights of linear layers by learnable univariate functions located at the nodes, yielding high expressivity with substantially fewer parameters. Finally, architectures such as the PINNsFormer [pinnsformer_2024] implement a transformer-based methodology for time-dependent data, translating each point $(t,r)$ into a short temporal sequence of $k$ steps that is fed into a standard encoder-decoder architecture, thus learning and enforcing temporal dependencies. In contrast, although PINNMamba [mamba_2025] shares certain similarities with the preceding approach, it replaces the MLP backbone with an encoder-only Mamba [ssm_2024] (selective state-space) network that models short, overlapping temporal subsequences. This design yields a continuous solution over the time of the physical problem, with consistency across overlaps enforced via a contrastive loss between the subsequences. All of these architectural approaches present distinct strengths and limitations, and each has been evaluated on a broad range of physics problems in their respective references. In our case, in Section 4, we undertake a comparative assessment of their performance in solving a variety of sub- and supercritical gravitational collapse problems of the scalar massless field minimally coupled to the spacetime metric.
@@
Regarding domain sampling, this step is also crucial to the methodology, since the collocation points—here the spacetime coordinates (t,r)(t,r)—serve as training data for the network. Several strategies for generating these points may be considered, ranging from simple, uniformly spaced grids or randomly generated samples for a particular probability distribution to more sophisticated techniques such as Sobol sequences [sobol_1967] or Latin hypercube sampling [stein_1987, mckay_2000]. Other approaches incorporating specialized deterministic distributions can also be considered. For example, methods that employ Chebyshev-distributed nodes in space combined with exponentially spaced temporal points enable more frequent and precise sampling near boundaries or localized features [vittorio_2025]. From a general perspective, any technique that involves non-uniform local refinement sampling can offer a strategy tailored to the specific problem, although such methods remain relatively uncommon in the PINN literature. Despite the fact that the performance of all these methods may vary depending on the specific PINN problem at hand, all of them share one common feature: they rely on fixed sampling strategies that remain static throughout the entire training process. In a comprehensive review, the authors of [florido_2024] concluded that there is no universal rule for defining a general and optimal sapling strategy for PINNs, saying that it is necessary to define the sampling in an intelligent manner and tailor the methodology to the particular scenario.
@@
Within the range of adaptive sampling methods, those based on probability distributions deserve special mention, highlighting [chenxi_2023] where the authors introduced the Residual-based Adaptive Distribution (RAD) which focused on resampling points with a probability proportional to their PDE residual. In studies such as [nabian_2021], the authors propose sampling techniques based on the significance assigned by the network to different regions of the training domain, finding that PINNs achieve faster convergence when the points are sampled according to a weighted probability density function (PDF) that depends on the loss. Some extensions of this idea are explored in works such as [song_2024], where a PDF is defined to continuously concentrate training points in regions with high loss values while still maintaining coverage in areas with lower errors. This balance is crucial, as completely neglecting low-error regions could lead to systematic forgetting by the PINN. Other studies focus more on repositioning the collocation points. For example, the RANG method (Residual-based Adaptive Node Generation) proposed in [peng_2022] treats the sampling points as variable-density radial basis function (RBF) nodes, allowing adaptive redistribution based on the residual.
3 Methodology
@@
In general, for the case of vanilla PINNs, the loss function employed, $\mathcal{L}$, given by Eq. (17) below, involves both the minimization of the governing PDEs, denoted by $\mathcal{L}_{\mathrm{PDE}}$, and the incorporation of the additional constraints, $\mathcal{L}_{i}$, among which the initial and boundary conditions are considered. A weight vector $\vec{\boldsymbol{\lambda}}=\left(\lambda_{\mathrm{PDE}},\lambda_{\mathcal{IC}},\lambda_{\mathcal{BC}},\ldots\right)$ is defined to combine the different components of the loss into a single objective function.
@@ 3.1 Construction of the loss function
Based on the definitions provided in the previous section and considering the compactness parameter $C$ defined in Eq. (13), the evolution equation for the scalar field can be expressed as,
@@
Spacetime tensor $(t,r)\in\mathbb{R}^{N_{t}N_{r}\times 2}$, residuals $\mathcal{L}(t,r)\in\mathbb{R}^{N_{t}N_{r}}$, adaptivity parameter $\Lambda>0$, domain bounds $r_{\min},r_{\max}]$.
@@
1. Reshape input: Reshape $r\to r_{\mathrm{old}}[n,i]$, $\mathcal{L}\to\mathcal{L}[n,i]$ of shape $N_{t}\times N_{r}$.

### v2

@@ Abstract
The gravitational collapse of a massless scalar field remains a demanding benchmark for numerical methods in numerical relativity, as it exhibits critical behavior at the boundary between dispersion and black hole formation. In this work we revisit this problem by relying on Physics-Informed Neural Networks (PINNs) as flexible solvers for partial differential equations, thereby providing a comparative assessment of several recent neural architectures. Building on the Einstein-massless-Klein-Gordon formulation in polar-areal coordinates, we consider four initial-value problems encompassing subcritical, critical, and supercritical regimes and use high-resolution finite-difference simulations as reference solutions. Our study is primarily comparative: we evaluate several state-of-the-art deep learning architectures, including vanilla and high-precision PINNs, sinusoidal-feature and quadratic-residual variants, and Kolmogorov-Arnold Networks, all trained under a common loss design that encodes the field equations, boundary conditions, and causal time-space enforcement, together with a novel adaptive spacetime sampling. Within this framework we also introduce ModPINN, a modest modification of standard PINNs that augments standard multilayer perceptrons with coordinate embeddings, quadratic layers, and other common ingredients in recent literature. This study shows that deep-learning-based methods can reproduce finite-difference solutions for the scalar field and the spacetime metric with competitive accuracy using significantly fewer collocation points than more traditional methodologies. While no single architecture dominates in all regimes, ModPINN achieves particularly stable and accurate solutions near criticality, indicating that suitably designed embeddings and adaptive sampling can enhance the robustness of PINNs for challenging gravitational-collapse scenarios.
@@ 1.1 Problem setup and basic equations
The study of gravitational collapse and black hole formation stands as one of the most interesting and revealing problems in modern physics. Beyond its astrophysical relevance, gravitational collapse provides an exceptional framework for testing the nonlinear, strong-field dynamics of general relativity (GR) and fundamental conjectures such as cosmic censorship. A particularly insightful and motivating aspect of this field arises in the so-called critical collapse problem, first discovered by Choptuik in the early 1990s [Choptuik_seminal], where the threshold between dispersion and black hole formation in the collapse of a massless scalar field was found to display a universal and self-similar behavior.
@@
The parameter $p$ corresponds to a specific characteristic of the initial distribution of the scalar field, $\phi(t=0,r):=\phi_{0}(r)$, where $t$ and $r$ denote the time and radial components, respectively, in polar-areal coordinates, $\{t,r,\varphi,\theta\}$. These coordinates help to express the time-dependent Schwarzschild-like metric
@@
The parameter $p$ naturally divides the entire set of initial conditions into two distinct categories: those that lead to the formation of a black hole during evolution and those that do not. Therefore, we can introduce the concept of a “threshold” when discussing the formation of an event horizon. Spacetimes evolving from initial conditions near this threshold exhibit properties that, from a mathematical perspective, are analogous to a critical phase transition in statistical mechanics (interested readers are referred to Sections 2.2 and 3.3 in [Gundlach2007]). In this manner, the parameter $p$ has a critical value $p^{*}$ separating “subcritical” solutions for which the mass dissipates ($p<p^{*}$) from “supercritical” solutions for which an event horizon forms ($p>p^{*}$), leading to the emergence of an apparent horizon.
@@
The set $(\Phi,\Pi)(t,r)$ in (8) is directly associated with the spatial and temporal derivatives of the scalar field, respectively. In terms of these variables, the set $(X,Y)(t,r)$ in (9) can be defined, leading to some relevant mathematical relations involving the derivatives of the mass aspect function, $m(t,r)$,
@@
For completeness, the spacetime curvature scalar $R$ can also be written in terms of this pair of variables as
@@
Within this formulation, this parameter satisfies $C\in[0,1)$, where the limit $C\to 1$ characterizes the onset of gravitational collapse. From a computational perspective, the formation of an apparent horizon in numerical simulations would be identified by imposing a threshold $\epsilon$ such that $1-C\leq\epsilon$. In this manner, classical numerical methods based on finite-difference methods (FDM) can effectively solve the field equations in both dispersive scenarios of $\phi(t,r)$ and cases involving gravitational collapse (e.g. see [Choptuik_1998]). Despite the coupling between the different variables, the dynamical evolution is governed by (5), while (6) and (7) serve as metric constraints.
@@ 1.2 Deep-learning-based methodologies
Despite the many numerical studies, it is worth noting that the most common methodology to solve the equations has been based on FDM. Only recent works have considered alternative methods to FDM to investigate gravitational collapse. In particular, the work of [finzi_2023] introduced a neural network-based solver that mitigates artificial neural network (ANN) ill-conditioning while exhibiting linear scalability with respect to the number of parameters. Other studies, such as [hatefi_2023], developed a model also based on ANNs to investigate critical phenomena in Einstein-axion-dilaton models, finding no evidence of self-similar behavior. On the other hand, [hatefi_2023_2] and [hatefi_2024] employ Bayesian deep learning models to study the axion-dilaton system, utilizing Sequential Monte Carlo (SMC) sampling and related algorithms to obtain direct estimates of the critical exponent $\gamma$.
All of these recent studies establish a foundation for the incorporation of deep learning (DL) techniques in the investigation of gravitational collapse scenarios. The present work proposes the adoption of novel methodologies based on ANNs tailored to solve systems of partial differential equations (PDEs) and extract their underlying physics. Since their initial appearance in the scientific literature [seminal_pinns], Physics-Informed Neural Networks (PINNs) have emerged as a robust alternative for solving numerical problems governed by PDEs. These methods leverage the universal approximation theorem of ANNs [Pinkus_1999] to minimize the governing physical equations, incorporating them into the loss function used to train the network parameters. This function is typically augmented with physical constraints—commonly referred to as “inductive biases”—such as initial and boundary conditions, additional physical information to be minimized, or even agreement with experimental data. Therefore, our primary objective will be to demonstrate that the PINN methodology is fully capable of solving the EMKG equations (5-7), effectively addressing both the field dispersion and the gravitational collapse scenarios. This approach is proposed as a viable alternative to traditional numerical methods.
The paper is structured as follows. Section 2 presents an overview of recent literature on PINNs, to motivate our study. Next, Section 3 introduces our PINN-based methodology, which consists of a modified version of the standard (vanilla) PINN; this section also details the specific formulation used to minimize the loss function. Section 4 presents the results obtained for various initial values of the scalar field, including the critical case, with particular emphasis on evaluating the predictive performance of the model in that regime. Finally, Section 5 presents our conclusions and outlines potential directions for future research.
@@ 2 Review of recent literature
In general terms, the choice of neural architecture is a critical factor, as the number of trainable parameters must be appropriate for the problem at hand, whatever the specific task may be—whether solving PDEs or addressing another application. Among the aspects that define a neural architecture, several may be highlighted: the activation functions employed; the number of layers (depth) and their size (width); whether residual or skip connections are incorporated between layers; and whether any form of embedding is used, among other considerations. Several studies [Yifan_2024] conclude, in a range of independent physical problems, that relatively shallow but wide networks often outperform deeper but narrower ones. This reveals a concern that usually appears throughout the literature on multilayer perceptrons (MLPs) applied to physics, motivating more recent developments such as Physics-Informed Residual Adaptive Networks (PirateNets) [piratenets_2024], which are introduced as residual architectures that allow the model to start from shallow networks, progressively “revealing” additional layers only when they are advantageous in reducing the objective cost function. These networks also employ embeddings of the physical coordinates, mapping them into a high-dimensional feature space using random Fourier features [Tancik_2020]. However, although these embeddings can mitigate the frequency bias typically observed in PINNs [Sifan_2021] and facilitate the solution of physical problems involving high-frequency behavior, the selection of the maximum frequency scale and the standard deviation of the Gaussian distribution used to sample the frequencies are critical; if chosen poorly, these become problem-dependent hyperparameters that can hinder performance. Other architectures, such as the sf-PINNs proposed in [Wong_2024], apply a mapping of the input coordinates to sinusoidal features, encoding them after the first layer, thus increasing the variability of the input gradient and also helping mitigate the said frequency bias.
@@
Stepping somewhat away from traditional architectures, methodological variants such as the Quadratic Residual Networks (QRes) [qres_2021] aim to replace the standard linear layer with an alternative that introduces second-order interactions between the trainable parameters and the input, becoming more parameter-efficient and faster to train, in general, on scientific problems. Within the most recent DL literature, Kolmogorov-Arnold Networks (KANs) [kan_2025] have emerged as an alternative to conventional MLPs, which have also been applied to PINNs. These mainly consist of replacing the trainable weights of linear layers with learnable univariate functions located at the nodes, yielding high expressivity with substantially fewer parameters. Finally, architectures such as the PINNsFormer [pinnsformer_2024] implement a transformer-based methodology for time-dependent data, translating each point $(t,r)$ into a short temporal sequence of $k$ steps that is fed into a standard encoder-decoder architecture, thus learning and enforcing temporal dependencies. In contrast, although PINNMamba [mamba_2025] shares certain similarities with the preceding approach, it replaces the MLP backbone with an encoder-only Mamba [ssm_2024] (selective state-space) network that models short, overlapping temporal subsequences. This design yields a continuous solution over the time of the physical problem, with consistency across overlaps enforced via a contrastive loss between the subsequences. All of these architectural approaches present distinct strengths and limitations, and each has been evaluated on a broad range of physics problems in their respective references. In our case, in Section 4, we undertake a comparative assessment of their performance in solving a variety of sub- and supercritical gravitational collapse problems of the massless scalar field minimally coupled to the spacetime metric.
@@
Regarding domain sampling, this step is also crucial to the methodology, since the collocation points—here the spacetime coordinates (t,r)(t,r)—serve as training data for the network. Several strategies for generating these points may be considered, ranging from simple, uniformly spaced grids or randomly generated samples for a particular probability distribution to more sophisticated techniques such as Sobol sequences [sobol_1967] or Latin hypercube sampling [stein_1987, mckay_2000]. Other approaches incorporating specialized deterministic distributions can also be considered. For example, methods that employ Chebyshev-distributed nodes in space combined with exponentially spaced temporal points enable more frequent and precise sampling near boundaries or localized features [vittorio_2025]. From a general perspective, any technique that involves non-uniform local refinement sampling can offer a strategy tailored to the specific problem, although such methods remain relatively uncommon in the PINN literature. Despite the fact that the performance of all these methods may vary depending on the specific PINN problem at hand, all of them share one common feature: they rely on fixed sampling strategies that remain static throughout the entire training process. In a comprehensive review, the authors of [florido_2024] concluded that there is no universal rule for defining a general and optimal sampling strategy for PINNs, saying that it is necessary to define the sampling in an intelligent manner and tailor the methodology to the particular scenario.
@@
Within the range of adaptive sampling methods, those based on probability distributions deserve special mention, highlighting [chenxi_2023] where the authors introduced the Residual-based Adaptive Distribution (RAD) which focused on resampling points with a probability proportional to their PDE residual. In studies such as [nabian_2021], the authors propose sampling techniques based on the significance assigned by the network to different regions of the training domain, finding that PINNs achieve faster convergence when the points are sampled according to a weighted probability density function (PDF) that depends on the loss. Some extensions of this idea are explored in works such as [Xiaotian_2024], where a PDF is defined to continuously concentrate training points in regions with high loss values while still maintaining coverage in areas with lower errors. This balance is crucial, as completely neglecting low-error regions could lead to systematic forgetting by the PINN. Other studies focus more on repositioning the collocation points. For example, the RANG method (Residual-based Adaptive Node Generation) proposed in [peng_2022] treats the sampling points as variable-density radial basis function (RBF) nodes, allowing adaptive redistribution based on the residual.
3 Methods
@@
In general, for the case of vanilla PINNs, the loss function employed, $\mathcal{L}$, given by (17) below, involves both the minimization of the governing PDEs, denoted by $\mathcal{L}_{\mathrm{PDE}}$, and the incorporation of the additional constraints, $\mathcal{L}_{i}$, among which the initial and boundary conditions are considered. A weight vector $\vec{\boldsymbol{\lambda}}=\left(\lambda_{\mathrm{PDE}},\lambda_{\mathcal{IC}},\lambda_{\mathcal{BC}},\ldots\right)$ is defined to combine the different components of the loss into a single objective function.
@@ 3.1 Construction of the loss function
Based on the definitions provided in the previous section and considering the compactness parameter $C$ defined in (13), the evolution equation for the scalar field can be expressed as,
@@
Spacetime tensor $(t,r)\in\mathbb{R}^{N_{\mathrm{t}}N_{\mathrm{r}}\times 2}$, residuals $\mathcal{L}(t,r)\in\mathbb{R}^{N_{\mathrm{t}}N_{\mathrm{r}}}$, adaptivity parameter $\Lambda>0$, domain bounds $r_{\min},r_{\max}]$.
@@
1. Reshape input: Reshape $r\to r_{\mathrm{old}}[n,i]$, $\mathcal{L}\to\mathcal{L}[n,i]$ of shape $N_{\mathrm{t}}\times N_{\mathrm{r}}$.